Divide 270 into three parts with a ratio of 2:3:4
Found 2 solutions by richwmiller, flame8855:
Answer by richwmiller(17219)

A=2X
B=3X
C=4X
A+B+C=270
2x+3x+4x=270
9x=270
x=30
A=60
B=90
C=120
check
60+90+120=270
270=270
ok

Answer by flame8855(424)